OpenStudy (anonymous):

PLEASE HELP!!! MEDALS!!!!!!! Medals! Seven subtracted from 3 times a number is equal to 9 more than 5 times the number. What is the number????? 3x-7=5x+9

OpenStudy (anonymous):

@poopsiedoodle @swissgirl @protected @mathmath333 @Algorithmic @anonymous_user @Teddyiswatshecallsme @KaceyMarie12345 @walkingdeadxx @arabpride @dejmccray Please Help me Last question!!

OpenStudy (arabpride):

3x-7=5x+9 -3x -3x -7 = 2x + 9 -9 -9 -16 = 2x /2 /2 -8 = x x = -8